OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 What is (f ◦ g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (jamesj):

By definition, (f ◦ g)(x) = f(g(x)) Now f(g(x)) = 5g(x) - 26, by definition of f(x), f(x) = 5x - 26 You complete it now by using the definition of g(x)

OpenStudy (jamesj):

That is f(g(x)) = 5g(x) - 26 = ... what?

OpenStudy (anonymous):

wouldn't you make it 5/x-26??

OpenStudy (jamesj):

No. f(x) = 5x - 26 hence, f(g(x)) = 5g(x) - 26 Now g(x) = 6x Hence f(g(x)) = 5g(x) - 26 = 5(6x) - 26, by definition of g(x) = ... what?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

sorry sorry 4 read it wrong